Show step by step:

Find the inverse for y=43sqaure root of x-2

1
Expert's answer
2021-03-23T06:24:11-0400

Solution:

"y=4\sqrt[3]{x-2}"

Interchange the variables:

"x=4\sqrt[3]{y-2}"

Solve for "y" :

"x=4\sqrt[3]{y-2}"

"x^3=4^3(y-2)"

"x^3=64(y-2)"

"y-2=\frac{x^3}{64}"

"y=\frac{x^3}{64}+2"

Replace the "y" with "f^{-1}(x)" to show the final answer:

"f^{-1}(x)=\frac{x^3}{64}+2"

Answer:

"f^{-1}(x)=\frac{x^3}{64}+2"